# Emergent spin and orbital angular momentum of light in twisted photonic bilayer

### v1

@@ Abstract
We demonstrate that the optical response of twisted photonic bilayers is sensitive to both spin angular momentum (SAM) and orbital angular momentum (OAM) of light. Due to the transfer of the angular momenta between the light and the bilayer, optical beams with zero SAM and OAM acquire SAM upon transmission and OAM upon reflection. Our numerical calculations and the developed analytical theory show that the values of the emergent SAM and OAM depend differently on the interlayer distance and can be controlled by tuning the twist angle. The predicted phenomena do not require light absorption and are caused by the helicity-dependent light diffraction on the moiré pattern, which inevitably occurs in the twisted structure, and SAM-OAM conversion. We also reveal strong SAM and OAM in the moiré-diffracted beams.
@@ I Introduction
In this paper, we show that twisted photonic bilayers exhibit both SAM- and OAM-dependent response. In particular, the incident optical beam with zero SAM (i.e., unpolarized) and zero OAM acquires SAM and OAM upon transmission and reflection. While the two phenomena might seem conceptually similar, we reveal that they have rather different nature, which leads to their very distinct properties: (i) SAM decays with the interlayer distance while OAM oscillates and (ii) SAM is observed in transmission while OAM is maximal in reflection. Moreover, we show that both phenomena do not require light absorption and are observed also in transparent bilayers. The emergence of SAM in the transmitted beam is caused by the diffraction of light by the moiré pattern, which inevitably occurs in the twisted structure and appears to be highly sensitive to the helicity of incident photons. We also study the parameters of the moiré diffracted beams and show that the beams possess significant SAM and OAM.
@@ II Model and Method
We consider a chiral bilayer consisting of two ($j=1,2$) thin dielectric slabs, Fig. 1. The layers are positioned at $z_{1}=0$, $z_{2}=d$ and rotated relatively to each other by the angle $\varphi$. The permittivity of each layer is assumed to be periodically modulated in the form of a square lattice with the period $2\pi/k$, so that the in-plane polarizations are given by
@@
where $\bm{\rho}$ is the in-plane coordinate, $\alpha_{0}$ is the uniform component of the layer polarizability and $\alpha_{1}$ is the amplitude of its spatial modulation, $\bm{k}_{j}$ and $\bm{k}^{\prime}_{j}$ are the reciprocal lattice vectors of the modulation, $|\bm{k}_{j}|=|\bm{k}^{\prime}_{j}|=k$, $\bm{k}_{j}\perp\bm{k}_{j}^{\prime}$, $\bm{k}_{1}\cdot\bm{k}_{2}=\bm{k}_{1}^{\prime}\cdot\bm{k}_{2}^{\prime}=k^{2}\cos\varphi$, and $\bm{E}_{\parallel}(\bm{\rho},z_{j})$ is the local in-plane electric field.
@@
with integer $n_{j}$ and $n_{j}^{\prime}$. We focus on the “metasurface” regime when the period of permittivity modulation in the layers is smaller than the wave length of incident light ($k>q_{0}$). Then, individual layers do not produce propagating diffracted beams. However, in twisted bilayers, because of coupling via near fields, the diffraction by the moiré pattern becomes possible, since its wave vector $K=|\bm{k}_{1}-\bm{k}_{2}|$ can be smaller than $q_{0}$, see Fig. 1.
@@
In the basis of circularly polarized plane waves ($\sigma=\pm 1$), the $D_{4}$ point-group symmetry of the bilayer, together with Lorentz reciprocity, imposes the constrains on the transmission and reflection coefficients at $\bm{q}_{\parallel}=0$. At transmission, the sign of the circular polarization is conserved, $t_{\sigma,\sigma^{\prime}}(0)=t_{\sigma}\delta_{\sigma,\sigma^{\prime}}$, however $t_{+}\neq t_{-}$ enabling the emergence of SAM $\langle s\rangle=(|t_{+}|^{2}-|t_{-}|^{2})/(|t_{+}|^{2}+|t_{-}|^{2})$ even for unpolarized incident beam. In reflection, $r_{\sigma,\sigma}(0)=0$ and $r_{+,-}(0)=r_{-,+}(0)\equiv r$ (we define the sign of the circular polarization with respect to the beam propagation direction), and the SAM vanishes. To understand the emergence of OAM one should consider optical beams constructed from plane waves with $\bm{q}_{\parallel}\neq 0$. At finite $\bm{q}_{\parallel}$, the above constrains on the transmission and reflection coefficients are violated. The SAM-OAM conversion processes such as $L_{z}=0,\;S_{z}\pm 1\rightarrow L_{z}=\pm 2,\;S_{z}\mp 1$, where $L_{z}$ and $S_{z}$ are the OAM and SAM projections, get allowed. The difference in the amplitudes of these processes results in the OAM in the reflected beam even if the incident beam carries no OAM.
Figure 2: (a) SAM $\langle s\rangle$ of the transmitted beam and (d) OAM $\langle l\rangle$ of the reflected beam as a function of the light frequency $\omega$ and the twist angle $\varphi$ for the interlayer distance $kd=1.34$. (b) SAM of the transmitted beam and (c) transmittance $T$, (e) OAM of the reflected beam and (f) reflectance $R$ as a function of the light frequency $\omega$ and the interlayer distance $d$ for the twist angle $\varphi=\pi/8$. Positions of the Fabry-Pérot resonances and the resonances associated with symmetric and antisymmetric waveguide modes are shown by solid, dashed, and dotted lines, respectively. The figures are calculated for $\alpha_{0}k=0.2$, $\alpha_{1}k=0.01$, $ak=50$, and $\gamma/(ck)=0.01$
@@ III Results
III.1 Transmitted and reflected beam
@@
First, we discuss the transmitted and reflected beams. We suppose that the incident beam is unpolarized, i.e., consists of incoherent $\sigma^{+}$ and $\sigma^{-}$ contributions. Figure 2 reveals that the transmitted beam acquires SAM [Figs. 2(a) and 2(b)] while the reflected beam acquires OAM [Figs. 2(d) and 2(e)]. Both effects are odd in the twist angle $\varphi$ and vanish at $\varphi=0,\pm\pi/4$ when the bilayer lacks chirality, see Figs. 2(a) and 2(d).
@@
The optical response of the bilayer contains the Fabry-Pérot resonances (solid lines in Fig. 2) and the resonances associated with the guided modes confined between the two layers (dashed and dotted lines in Fig. 2[37]. The frequencies of the symmetric and antisymmetric guided modes with the in-plane wave vector $\bm{g}$ and of the Fabry-Pérot resonances with $\bm{g}=0$ in the limit of vanishing $\alpha_{1}$ are determined by
$2\pi q_{0}^{2}\alpha_{0}\left(1\pm{\rm e}^{-\sqrt{g^{2}-q_{0}^{2}}d}\right)=\sqrt{g^{2}-q_{0}^{2}}\,,$ (4)
where $\bm{g}$ are the diffraction vectors determined by Eq. (3). The in-plane modulation of the permittivity $\alpha_{1}$ couples the guided modes to the free-space modes, therefore they acquire finite lifetime and can be excited by normally incident beam. The resonances are clearly seen in the transmittance and reflectance, see Figs. 2(c) and 2(f), respectively. Both SAM and OAM are enhanced at the resonances. For our case of weak dielectric modulation $\alpha_{1}k\ll 1$, the enhancement is particularly pronounced at the frequencies of the guided modes with the first-order diffraction wave vectors $|\bm{g}|=k$ (black dashed and dotted lines in Fig. 2). The OAM of the reflected beam is additionally increased at the intersections of the guided modes with the Fabry-Pérot modes, where the conventional reflection is suppressed.
@@
Though the SAM (circular polarization) of the beam transmitted through the chiral bilayer may seem to be natural, its microscopic origin is in fact unusual and differs from the known mechanism. In conventional chiral media, the circular polarization of the transmitted (initially unpolarized) beam is caused by the circular dichroism, i.e., the difference in the absorption of $\sigma^{+}$ and $\sigma^{-}$ waves. Here, the circular polarization emerges even in the absence of absorption. Surprisingly, the transmission coefficients for $\sigma^{+}$ and $\sigma^{-}$ photons are different although the photons retain their helicity and the reflection coefficients for $\sigma^{+}$ and $\sigma^{-}$ photons coincide. This puzzle is solved by taking into account the moiré pattern of twisted bilayers and unavoidable diffraction by the moiré wave vectors. In chiral bilayers, the intensity of the diffracted beams depends on the circular polarization of the incident beam. This difference in the decay of incident $\sigma^{+}$ and $\sigma^{-}$ photons over the diffraction channels gives rise to the circular polarization of the transmitted beam.
@@
Simple analytic expressions for the SAM and OAM can be obtained in the case of small layer polarizability, $\alpha_{0}k,\alpha_{1}k\ll 1$, i.e. away from the resonances, see Appendix B. The spin-dependent contribution to the transmission coefficient comes from the fourth-order diffraction processes, e.g., consequently by $\bm{k}_{1}$, $-\bm{k}_{2}$, $-\bm{k}_{1}$, $\bm{k}_{2}$. For small $\varphi$, the SAM of the transmitted beam assumes the form
@@
where $\varkappa=\sqrt{k^{2}-q_{0}^{2}}$ is the reciprocal decay length of the near field. As the function of the interlayer distance $d$, the SAM rises linearly at small $d$, reaches maximum at $d\sim 1/\varkappa$, and then decreases exponentially at large $d$. Such a non-monotonous dependence also persists in the full calculation Fig. 2(b), where the SAM is shown as a function of the light frequency and the interlayer distance. The exponential decrease at large $d$ clearly indicates that the SAM originates from the near field interaction between the layers. Indeed, at large interlayer distances $d\gg 1/\varkappa$, the interaction of light with each layer can be considered independently. Since the individual layer possesses high symmetry ($D_{4h}$ point group), it exhibits polarization-independent transmission and reflection in the paraxial approximation, so that the SAM vanishes.
The OAM for the reflected beam arises already in the second-order diffraction processes, e.g., consequently by $\bm{k}_{1}$ and $-\bm{k}_{1}$. Such processes occur from individual layers and lead to distortions of the Gaussian beam shape. Interference of the distortions induced by the first and the second layer leads to the OAM
@@
Similarly to the SAM, the OAM vanishes at $\varphi=0$ and reaches its maxima at the twist angle $\varphi=\pm\pi/8$. However, being determined by the interference of the far fields rather than by the near fields, the OAM does not decay with the increase of the interlayer distance. In transmission, the optical paths of the beams distorted by the first and second layer are the same, suppressing their interference and leading to a much smaller value of the OAM than for the reflected beam.
III.2 Moiré diffracted beams

### v2

@@ Abstract
We demonstrate that the optical response of twisted photonic bilayers, photonic counterparts of van der Waals structures, is sensitive to both spin angular momentum (SAM) and orbital angular momentum (OAM) of light. A beam of unpolarized light with zero angular momentum acquires SAM in transmission and OAM in reflection. The developed analytical theory and numerical calculations show that the SAM and OAM arise from distinct microscopic mechanisms and depend differently on the interlayer distance. The predicted phenomena do not require light absorption and are caused by the photon-helicity-dependent light diffraction by the moiré pattern, which inevitably occurs in the twisted structure, and the SAM-OAM conversion processes. We also reveal strong SAM and OAM in the moiré-diffracted beams. Our findings uncover a profound connection between the emergent SAM and OAM in twisted photonic systems offering new possibilities for angular-momentum-resolved light-matter interactions.
@@ I Introduction
In this paper, we show that twisted photonic bilayers exhibit both SAM- and OAM-dependent response. In particular, when the bilayer is illuminated by a beam of unpolarized light with zero angular momentum, the transmitted and reflected beams acquire SAM and OAM. While the two phenomena have common phenomenological origin in the chirality of the system, we reveal that they have different microscopic nature, which leads to their very distinct properties. (i) SAM originates from the near-field interaction between the layers and decays with the interlayer distance while OAM comes from the interference of the SAM-OAM conversion in the layers and oscillates with the interlayer distance with no decay. (ii) SAM is observed in transmission while OAM is maximal in reflection. Moreover, we show that both phenomena do not require light absorption and are observed also in transparent bilayers. The emergence of SAM in the transmitted beam is caused by the diffraction of light by the moiré pattern, which inevitably occurs in the twisted structure and appears to be highly sensitive to the helicity of incident photons. We also study the parameters of the moiré diffracted beams and show that the beams possess significant SAM and OAM.
@@ II Model and Method
We consider a chiral bilayer consisting of two ($j=1,2$) thin dielectric slabs, Fig. 1. The layers are positioned at $z_{1}=0$, $z_{2}=d$ and rotated relatively to each other by the angle $\varphi$. Each layer is characterized by the 2D permittivity which is periodically modulated in the form of a square lattice with the period $2\pi/k$, so that the in-plane 2D polarizations are given by
@@
where $\bm{\rho}$ is the in-plane coordinate, $\alpha_{0}$ is the uniform component of the layer polarizability and $\alpha_{1}$ is the amplitude of its spatial modulation, $\bm{k}_{j}$ and $\bm{k}^{\prime}_{j}$ are the reciprocal lattice vectors of the modulation, $|\bm{k}_{j}|=|\bm{k}^{\prime}_{j}|=k$, $\bm{k}_{j}\perp\bm{k}_{j}^{\prime}$, $\bm{k}_{1}\cdot\bm{k}_{2}=\bm{k}_{1}^{\prime}\cdot\bm{k}_{2}^{\prime}=k^{2}\cos\varphi$, and $\bm{E}_{\parallel}(\bm{\rho},z_{j})$ is the local in-plane electric field. The 2D polarizability depends on the real thickness of the dielectric slab.
@@
with integer $n_{j}$ and $n_{j}^{\prime}$. We focus on the “metasurface” regime when the period of permittivity modulation in the layers is smaller than the wave length of incident light ($k>q_{0}$). Then, individual layers do not produce propagating diffracted beams. However, in twisted bilayers, because of coupling via near fields, the diffraction by the moiré pattern becomes possible, since its wave vector $K=|\bm{k}_{1}-\bm{k}_{2}|$ can be smaller than $q_{0}$, see Fig. 1. In numerical solution, to smoothen the sharp resonances we use the regularization $\omega\to\omega+{\rm i}\gamma$ with small $\gamma$ which corresponds to the medium with weak losses.
@@
Figure 2: (a) SAM $\langle s\rangle$ of the transmitted beam and (d) OAM $\langle l\rangle$ of the reflected beam as a function of the light frequency $\omega$ and the twist angle $\varphi$ for the interlayer distance $kd=1.34$. (b) SAM of the transmitted beam and (c) transmittance $T$, (e) OAM of the reflected beam and (f) reflectance $R$ as a function of the light frequency $\omega$ and the interlayer distance $d$ for the twist angle $\varphi=\pi/8$. Positions of the Fabry-Pérot resonances and the resonances associated with symmetric and antisymmetric waveguide modes are shown by solid, dashed, and dotted lines, respectively. The figures are calculated for $\alpha_{0}k=0.2$, $\alpha_{1}k=0.01$, $ak=50$, and $\gamma/(ck)=0.01$.
@@ III Results
III.1 General notes
Before presenting the results of calculations, we draw some general conclusions based on symmetry arguments. First, we note that individual photonic layers constituting the bilayer are described by the $D_{4h}$ point group, which is achiral and centrosymmetric. The transmission and reflection coefficients of individual layers are polarization independent for normally incident beams. The bilayer becomes chiral if the layers are stacked with the twist providing the twist angle $\varphi\neq\pi n/4$, where $n$ is integer, and the interlayer distance $d\neq 0$. The symmetry of such a bilayer is reduced to the $D_{4}$ point group, which enables the emergence of SAM and OAM for normally incident beams.
Second, one can formulate optimal conditions on the interlayer distance $d$. It should not be too small since the bilayer lacks chirality at $d=0$. On the other hand, it should not be too large otherwise the layers get optically decoupled. The interaction via near fields, essential for the emergence of SAM and moiré diffraction, is effective if the interlayer distance $d$ is comparable to the wavelength of permittivity modulation $2\pi/k$. Assuming that $k$ is of the order of the photon wave vector $q_{0}$ we conclude that the optimal values of $d$ are in the $\mu$m scale for the visible spectral range and in the sub-mm scale for terahertz radiation.
Third, the $D_{4}$ point-group symmetry of the chiral bilayer together with the Lorentz reciprocity imposes the constraints on the transmission and reflection coefficients at $\bm{q}_{\parallel}=0$. We formulate them in the basis of circularly polarized plane waves ($\sigma=\pm 1$). At transmission, the sign of the circular polarization is conserved, $t_{\sigma,\sigma^{\prime}}(0)=t_{\sigma}\delta_{\sigma,\sigma^{\prime}}$. However, the coefficients $t_{+}$ and $t_{-}$ can be different, which enables the emergence of SAM $\langle s\rangle=(|t_{+}|^{2}-|t_{-}|^{2})/(|t_{+}|^{2}+|t_{-}|^{2})$ even for unpolarized incident beam. In reflection, $r_{\sigma,\sigma}(0)=0$ and $r_{+,-}(0)=r_{-,+}(0)\equiv r$ (we define the sign of the circular polarization with respect to the beam propagation direction), and the SAM $\langle s\rangle=(|r_{+}|^{2}-|r_{-}|^{2})/(|r_{+}|^{2}+|r_{-}|^{2})$ vanishes. Thus, the transmitted light acquires SAM while the reflected light does not.
To understand the emergence of OAM one should consider optical beams constructed from the plane waves with nonzero $\bm{q}_{\parallel}$. At finite $\bm{q}_{\parallel}$, the above constraints on the transmission and reflection coefficients are violated. The SAM-OAM conversion processes take place such as $L_{z}=0,\;S_{z}=\pm 1\rightarrow L_{z}=\pm 2,\;S_{z}=\mp 1$ and $L_{z}=0,\;S_{z}=\pm 1\rightarrow L_{z}=\mp 2,\;S_{z}=\mp 1$, where $L_{z}$ and $S_{z}$ are the OAM and SAM projections on the incident beam axis. The former processes conserve the total photon angular momentum projection $L_{z}+S_{z}$. In the latter processes, the total photon angular momentum projection changes by ±$\pm 4$ and the difference is converted into the mechanical rotation of the twisted bilayer. They are allowed by 4-fold rotation symmetry of the bilayer and play a major role in the OAM emergence. The difference in the rates of the conversion processes results in the OAM in the reflected beam even if the incident beam carries no OAM ($L_{z}=0$) and consists of the equal portions of the photons with $S_{z}=\pm 1$.
III.2 Transmitted and reflected beam
@@
First, we discuss the transmitted and reflected beams. We suppose that the incident beam is unpolarized, i.e., consists of incoherent $\sigma^{+}$ and $\sigma^{-}$ contributions. Figure 2 reveals that the transmitted beam acquires SAM [Figs. 2(a) and 2(b)] while the reflected beam acquires OAM [Figs. 2(d) and 2(e)]. Both effects are odd in the twist angle $\varphi$ and vanish at $\varphi=0,\pm\pi/4$ when the bilayer lacks chirality, see Figs. 2(a) and 2(d). The obtained values of the SAM $\langle s\rangle$ and the OAM $\langle l\rangle$ are non-integer. It means that the beams are the mixtures of the states with integer SAM and OAM projections with different portions. In our case, the mixture is formed by the states with $S_{z}=\pm 1$ and $L_{z}=0,\pm 2$.
@@
The optical response of the bilayer contains the Fabry-Pérot resonances (solid lines in Fig. 2) and the resonances associated with the guided modes confined between the two layers (dashed and dotted lines in Fig. 2[37]. The resonances are clearly seen in the transmittance $T$ and reflectance $R$, see Figs. 2(c) and 2(f), respectively. The Fabry-Pérot interference leads to the oscillations in the transmittance and reflectance almost from 0 to 1 and is well described by the equations
$T=\frac{1}{|(1-2\pi{\rm i}\alpha_{0}q_{0})^{2}+(2\pi\alpha_{0}q_{0})^{2}{\rm e}^{2{\rm i}q_{0}d}|^{2}},\;\;R=1-T,$ (4)
which are derived neglecting the lateral modulation of the permittivity. The resonances associated with the guided modes are much weaker in the transmittance and reflectance. They get visible because the in-plane modulation of the permittivity $\alpha_{1}$ couples the guided modes to the free-space modes. The guided modes acquire finite lifetime and can be excited by normally incident beam. The frequencies of the symmetric and antisymmetric guided modes with the in-plane wave vector $\bm{g}$ and the Fabry-Pérot resonances with $\bm{g}=0$ in the limit of small $\alpha_{1}$ are determined by the equation
$2\pi q_{0}^{2}\alpha_{0}\left(1\pm{\rm e}^{-\sqrt{g^{2}-q_{0}^{2}}d}\right)=\sqrt{g^{2}-q_{0}^{2}}\,.$ (5)
Both SAM and OAM are enhanced at the resonances. For our case of weak dielectric modulation $\alpha_{1}k\ll 1$, the enhancement is particularly pronounced at the frequencies of the guided modes with the first-order diffraction wave vectors $|\bm{g}|=k$ (black dashed and dotted lines in Fig. 2). At the resonances, the SAM changes sign as the function of the light frequency. This behavior is similar to the spectral dependence of the circular dichroism at the exciton resonance in bulk gyrotropic crystals [2], quantum wells [39], and chiral stacks of 2D semiconductors [36]. Microscopically, the inversion originates from the splitting of the spectral positions of the resonances which are excited by $\sigma^{+}$ and $\sigma^{-}$ circularly polarized waves with the given wave vector. The OAM of the reflected beam is additionally increased at the intersections of the guided modes with the Fabry-Pérot modes, where the conventional reflection is suppressed.
Figure 3: (a) SAM $\langle s\rangle$ of the transmitted beam and (b) bilayer transmittance $T$ as a function on the light frequency in the vicinity of a guided mode resonance. The curves are calculated for $kd=2$, $\varphi=\pi/8$, and three different rates of energy losses $\gamma$. Other parameters are listed in the caption to Fig. 2.
The amplitudes and widths of the guided mode resonances in Fig. 2 are determined by the background energy losses which are included in the numerical calculations. The used rate of the losses $\gamma=10^{-2}ck$ is small but still exceeds the radiative broadening of the resonances determined by their coupling to propagating modes. Figure 3 shows the evolution of the SAM and the bilayer transmittance at a guided mode resonance with the decrease of the losses. The resonance gets much narrower and the SAM values reach ±$\pm 1$. Further decrease of $\gamma$ does not affect much the results indicating that the limit of radiative broadening is almost achieved. Figure 3 reveals that the fine tuning of the guided mode resonances to the desired light frequency enables the achievement of high SAM in the transmitted light even in a single twisted bilayer.
@@
Though the SAM (circular polarization) of the beam transmitted through the chiral bilayer may seem to be natural, its microscopic origin is in fact unusual and differs from the known mechanism. In conventional chiral media, the circular polarization of the transmitted (initially unpolarized) beam is caused by the circular dichroism, i.e., the difference in the absorption of $\sigma^{+}$ and $\sigma^{-}$ waves. Here, the circular polarization emerges even in the absence of losses, which is particularly pronounced in Fig. 3. Surprisingly, the transmission coefficients for $\sigma^{+}$ and $\sigma^{-}$ photons are different although there is no absorption, the photons retain their helicity and the reflection coefficients for $\sigma^{+}$ and $\sigma^{-}$ photons coincide. This puzzle is solved by taking into account the channels of photon decay associated with the moiré diffraction. Indeed, twisting the bilayer produces the moiré pattern and unavoidable diffraction of light by the moiré wave vectors. In chiral bilayers, the intensity of the diffracted beams depends on the circular polarization of the incident beam. This difference in the decay of incident $\sigma^{+}$ and $\sigma^{-}$ photons over the moiré diffraction channels gives rise to the circular polarization of the transmitted beam.
@@
Simple analytic expressions for the SAM and OAM can be obtained in the perturbation theory in the case of small layer polarizability, $\alpha_{0}k,\alpha_{1}k\ll 1$, see Appendix B. The spin-dependent contribution to the transmission coefficient comes from the fourth-order diffraction processes, e.g., consequently by $\bm{k}_{1}$, $-\bm{k}_{2}$, $-\bm{k}_{1}$, $\bm{k}_{2}$. For small $\varphi$, the SAM of the transmitted beam assumes the form
@@
where $\varkappa=\sqrt{k^{2}-q_{0}^{2}}$ is the reciprocal decay length of the near field. As the function of the interlayer distance $d$, the SAM rises linearly at small $d$, reaches maximum at $d\sim 1/\varkappa$, and oscillates and decreases exponentially at large $d$. Such a non-monotonous dependence also persists in the full calculation Fig. 2(b), where the SAM is shown as a function of the light frequency and the interlayer distance. The oscillations originate from the Fabry-Pérot interference in the bilayer. The exponential decrease at large $d$ clearly indicates that the SAM originates from the near-field coupling of the layers. Indeed, at large interlayer distances $d\gg 1/\varkappa$, the interaction of light with each layer can be considered independently. Since the individual layer possesses high symmetry ($D_{4h}$ point group), it exhibits polarization-independent transmission and reflection in the paraxial approximation, so that the SAM vanishes.
The OAM for the reflected beam arises already in the second-order diffraction processes, e.g., consequently by $\bm{k}_{1}$ and $-\bm{k}_{1}$. Such processes occur from individual layers and lead to distortions of the phase surface of the Gaussian beam. Interference of the distortions induced by the first and the second layers leads to the OAM
@@
Figure 4: Spatial distribution of SAM in the plane $z=d/2$ between the layers for the twist angle (a) $\varphi=\pi/8$ and (b) $\varphi=\pi/15$. The figures are calculated for $kd=2$, $\omega/(ck)=0.816$, and unpolarized incident light. Other parameters are listed in the caption to Fig. 2.
Similarly to the SAM, the OAM vanishes at $\varphi=0,\pm\pi/4$ and reaches its maxima at the twist angle $\varphi=\pm\pi/8$. The reflected beam OAM also oscillates as the function of the interlayer distance $d$, which originates from the Fabry-Pérot interference. However, being determined by the interference of the far fields rather than by the near fields, the OAM does not decay with the increase of the interlayer distance. In transmission, the optical paths of the beams distorted by the first and second layers are the same, suppressing their interference and leading to a much smaller value of the OAM than for the reflected beam.
In conclusion of this subsection, we comment that the near field in a twisted bilayer has a complex structure being the superposition of propagating modes and evanescence modes with different wave vectors $\bm{g}$. Figure 4 shows the spatial distributions of the SAM in the plane $z=d/2$ centered between the layers. The distributions are calculated for the parameters $kd=2$ and $\omega/(ck)=0.816$ corresponding to a guided mode resonance, see Fig. 2(b), and two different twist angles $\varphi$. As expected, the SAM oscillates in the plane. The local values of the near field SAM are much higher than the SAM in the transmitted light, cf. Fig. 2 and Fig. 4.
III.3 Moiré diffracted beams